Write the equation of the line that passes through (-2,5) and (7,10)
Answer by checkley75(3666)

first we find the slope (m)
(y2-y1)/(x2-x1)
(10-5)/(7+2)
5/9 is the slope.
now substitute one set of (x,y) values in the equation Y=MmX+b & solve for b
5=5/9*-2+b
5=-10/9+b
b=5+10/9
b=(45+10)/9
b=55/9
thus this line equation is
y=5/9x+55/9